OpenStudy (anonymous):

Plzz help Simplify cube root of 7 over fifth root of 7.

OpenStudy (anonymous):

A. 7 to the power of 1 over 5 B. 7 to the power of 8 over 15 C. 7 to the power of 5 over 3 D. 7 to the power of 2 over 15

OpenStudy (anonymous):

@They_Call_Me_Narii can you help me plz :)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

is that right bro? ^ @BoraSemiz

OpenStudy (anonymous):

I thkin what you want to solve this: \[\sqrt[3]{\frac{ 7 }{5}} = ? \sqrt[7]{\frac{ x }{ y }}\] right ? (You are going to find x and y, right ?)

OpenStudy (anonymous):

yea ^

OpenStudy (anonymous):

Medal Me

OpenStudy (anonymous):

@They_Call_Me_Narii could you explain :)

OpenStudy (anonymous):

Cube Root Means 3 And Fifth Root Is Five And 5+3=8 So Its 7 to the power of 8 over 15

OpenStudy (anonymous):

k thx @They_Call_Me_Narii